# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v2

@@ 1 Introduction
Mixed systems of bosons and fermions are at the base of our understanding of the physical world. They feature in the Standard Model of elementary-particle physics as gauge and matter fields, respectively, in high-density quark matter as diquarks interacting with unpaired quarks [1], in nuclear matter in the interacting boson-fermion model for atomic nuclei with an odd number of protons or neutrons [2], and as boson-mediated interactions in soft [3] as well as condensed [4] matter. Ultra-cold atomic gases provide an exceptionally versatile platform for the simulation of quantum matter [5]. In the case of Bose-Fermi (BF) atomic mixtures, several realizations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have permitted to explore a wide range of phenomena like phase separation [22], polarons [13, 17, 23, 26, 31], dual superfluidity [15, 16, 19, 20], collective excitations [34, 39], mediated interactions [28, 35, 36, 37], and Feshbach molecules [6, 8, 10, 25, 30, 33].
@@
In a parallel fashion, BF mixtures have also been realized in semiconductor nanostructures thanks to the recent advent of atomically thin transition metal dichalcogenides (TMD) [40, 41, 42, 43, 44, 45]. These materials intrinsically offer a remarkable opportunity to investigate the physics in two dimensions (2D) where topological phases play a crucial role [46, 47]. The so-called spin-valley locking, originating from the strong spin-orbit coupling of the transition metal atoms, allows for a complete distinguishability between electrons (and/or holes) populating the two spin-polarized valleys at Brillouin points $K$ and $K^{\prime}$ [48]. Exciton formation and charge doping are therefore two distinct processes affecting separate assemblies of electrons (and/or holes). In this scenario, the production of a BF mixture in a two-dimensional solid-state setting becomes therefore possible, with bosons corresponding to tightly bound excitons and fermions to doped charges. The recent detection of electrically tunable two-dimensional Feshbach resonances in TMD bilayers has put this target on more solid ground [43, 49].
@@
However, the unique flexibility of atomic gases is unrivaled when the need arises to manipulate dimensionality and interaction strength at the same time. Confinement-induced resonances accomplish this purpose by allowing for the realization of a dimensional crossover, starting from 3D confinement down to (quasi-)2D one. If the energy level spacing of the 3D trapping system (or equivalently the confinement length) is kept much larger than the energy (shorter than the length) scales of the intervening scattering processes, the system becomes effectively 2D and the ensuing 2D scattering length can be manipulated by varying either the scattering length associated with the original 3D Feshbach resonance and/or the confinement length, yet ensuring the quasi-2D condition above remains satisfied [50, 51, 52]. This situation is particularly apt for BF mixtures, for which one would need to adjust independently the BF and the boson-boson (BB) interactions owing to the mechanical stability of the assembly being conditioned primarily by that of the Bose component [53]. Having at one’s disposal two independent degrees of freedom in the 3D scattering length and the confinement length, one can then correspondingly tune the effective BF and BB interactions (even though in a highly non-linear fashion [50]) and simultaneously analyze the stability of different phases of the mixture across the whole resonant regime.
@@
Under the above scenario, it is crucial to explore BF mixtures with a tunable BF attraction in a 2D setting. Until now, theoretical efforts have concentrated mainly on 3D settings [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. In 2D, theoretical work has been sparse so far [87, 88, 89, 90] and has primarily dealt with phase competition in optical lattices [89, 87], phase stability in atomic traps [88], or with the bosonic impurity limit of a boson density ${n_{\mathrm{B}}}$ much smaller than the fermion density ${n_{\mathrm{F}}}$ [90].
In this work, we study a 2D BF mixture in homogeneous space with boson and fermion densities spanning from the bosonic impurity limit (${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$) to the matched density regime (${n_{\mathrm{B}}}={n_{\mathrm{F}}}$). Similarly to the analyses carried out in 3D [76, 80, 84, 83, 86, 91], we are here interested in the competition between boson condensation and BF pairing in mixtures with an attractive and tunable BF interaction at zero temperature. For this reason, we consider concentrations of bosons $x={n_{\mathrm{B}}}/{n_{\mathrm{F}}}\leq 1$, such that a full competition between pairing and condensation is allowed. The BB interaction (when considered) is taken repulsive for mechanical stability reasons. Numerical calculations focus on the case of equal boson and fermion masses in order to reduce the computational effort. However, to facilitate possible future studies, the theoretical formalism is developed for generic boson and fermion masses $m_{\rm B}$ and $m_{\rm F}$.
Our main results are as follows. (i) We find that the condensate is progressively depleted as the BF attraction increases, in analogy with the 3D case. However, while in 3D the condensate vanishes beyond a critical coupling strength, in 2D the condensate does not exactly vanish but rather becomes exponentially small with the coupling strength. (ii) Similarly to the 3D case, we uncover a nearly universal behavior of the condensate fraction, the bosonic momentum distribution, and Tan’s contact parameter with respect to the boson concentration. (iii) BF pairs form for sufficiently strong BF attraction. Quite generally, these composite fermions are hybridized states between a pure molecular bound state and an unpaired state made of a fermionic atom and a boson belonging to the condensate. We find that these hybridized states produce quasi-particles with rather peculiar Fermi liquid features. (iv) The boson momentum distribution function shows quite a rich behavior when varying the BF coupling strength and boson concentration, which could in principle be tested in future experiments with BF mixtures.
@@ 2.1 The system Hamiltonian
On physical grounds, one expects the following picture when the BF coupling strength is varied. In the weak-coupling limit, bosons are expected to be almost fully condensed, with a small depletion of the condensate density $n_{0}$ determined just by the weak BB repulsion [94] for $g\to-\infty$. Fermions instead fill up a Fermi sphere. In the strong-coupling limit, bosons are instead expected to be expelled out of the condensate and form molecular fermions that fill up a Fermi sphere. We will discuss below to what extent, within the present approach, this simple picture continues to hold also in two dimensions.
@@ 2.2 Many-body $T$-matrix
A warning is however in order. In the limiting case of a single impurity in a Fermi gas, it has been shown that the present approach misses the polaron-to-molecule transition in 2D [97], which is instead found by more sophisticated approaches [98, 99]. We should thus bear in mind that such a shortcoming of the present approximation, which is due to an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory, may extend also at finite boson concentration $x$, in particular in the limit $x\to 0$ that approaches the single impurity case.
@@
Let us see now the rationale behind the choice of diagrams made in [91], which we wish to extend to 2D. Since the pioneering work by Nozières and Schmitt-Rink [100] on the related problem of the BCS-BEC crossover in two-component Fermi gases, it is well-known that ladder diagrams are able to capture pairing (molecular) correlations in the normal phase (see also [101] for a review). It is only after the inclusion of this class of diagrams that the superfluid critical temperature recovers the Bose-Einstein condensation temperature in the strong-coupling limit of the BCS-BEC crossover. In addition, for the same contact potential we are considering, ladder diagrams provide the leading self-energy in the weak-coupling limit of the BCS-BEC crossover [102]. They thus provide a sensible scheme to describe the whole BCS-BEC crossover, even in the intermediate coupling region in which fully controlled approximations are not available.
@@
whereby the $\Theta$ functions appearing in the integrand correspond to the Fermi and Bose distributions in the zero temperature limit. In particular, for the concentrations $x\leq 1$ and the range of BF and BB coupling strengths considered in this work, ${\mu_{\mathrm{B}}}$ is always negative, hence $\Theta(-\xi_{\mathbf{k}}^{\mathrm{B}})=0$.
@@ 2.3 Boson and fermion self-energies
The bosonic self-energy is made of two terms: originating, respectively, from BF and BB interactions. For the former, we proceed like in the corresponding problem in the BCS-BEC crossover: the self-energy of one species is obtained by closing the $T$-matrix by a propagator of the other species.
@@ 2.4 Green’s functions, momentum distributions, and densities
Finally, due to the presence of a bosonic component at zero temperature, the Hugenholtz-Pines condition ([107]) for the corresponding chemical potential is imposed as
@@ 3.2 Fermionic self-energy and momentum distribution
The frequency integration in Eq. (12) is computed via a contour integration on the left-hand side of the complex plane, owing to the presence of the convergence factor $e^{i\omega 0^{+}}$. The pole originating from $G_{\mathrm{B}}^{0}({\mathbf{P}}-{\mathbf{k}},\Omega-\omega)$ is located on the right-hand side of the complex plane because $\mu_{\rm B}<0$, and does not contribute to the integral. The contributions from the poles of $T_{\mathrm{SC}}$ yield instead
@@ 3.4 Complete system of semi-analytic equations in closed form and their solutions
The momentum distribution of the composite fermions $n_{\mathrm{CF}}(\mathrm{P})=v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is displayed in Fig. 5 for a number of concentrations corresponding to those in Fig. 4 along with the case $x=1$. In the latter case, the mixture is essentially made of a single component gas of non-interacting point-like fermionic molecules, as illustrated by the Fermi step behavior with occupation number approaching unity. As $x$ decreases, interaction effects, originating from the hybridization of the molecular and atomic states, become significant and generate a rather peculiar behavior of the momentum distribution function, as evidenced by its upward bending with increasing momentum. (A similar behavior was also found in [71] for a BF mixture in a 1D lattice.) Moreover, we notice that the Fermi step, in particular for small concentrations, is located well above the expected position, if one would assume the validity of the Luttinger theorem for the composite fermions.
@@
The Luttinger theorem states that the volume of the Fermi sphere of a Fermi liquid does not depend on interaction, and thus coincides with the volume of the Fermi sphere of the corresponding non-interacting Fermi gas [113]. Its validity has been extended to density-imbalanced two-component Fermi gases [114] and, within a two-channel model, to BF mixtures [65]. In both cases, it was shown that, provided the system is in the normal phase, the volumes of the Fermi spheres are separately conserved for the two Fermi components (in the BF mixture case, the two components correspond to renormalized fermionic atoms and molecules). This is indeed what is found in the strong-coupling limit of a 3D BF mixture in [84], as in this limit the condensate vanishes identically. In contrast, in the present 2D case the condensate density never identically vanishes, and in particular the hybridization energy $\Delta_{0}$ remains finite even for $g\to\infty$ (except in the trivial case $x=0$ and in the more interesting case $x=1$).
@@
In a BF mixture with a condensate, it was shown in [65] that it is the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules that should be conserved. Fig. 5 shows a clear violation of the latter version of the Luttinger theorem since the molecular Fermi momentum is always larger than $k_{\mathrm{F}}$, which is the Fermi momentum that should determine the total volume (i.e the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules). We ascribe this shortcoming to two factors. First, the large values of the hybridization energy-scale $\Delta_{0}$ obtained in our theory, which are in turn caused by an overestimation of the atom-molecule repulsion (see Sec. 3.1), eventually lead to an overpopulation of the condensate. So, the simple strong-coupling picture of a Fermi-Fermi mixture of unpaired fermions and molecules obeying the Luttinger theorem is never reached in 2D. Second, as stressed in [114] and explicitly shown in [115], when dealing with approximate theories the validity of the Luttinger theorem is not guaranteed for schemes, like the present one, that do not use fully self-consistent fermionic Green’s functions in their formulation.
@@
Nevertheless, the breakdown of the Luttinger theorem in 2D for the hybridized composite fermions is compensated, as far as the integrated quantity $n_{\mathrm{CF}}=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}n_{\rm CF}({\mathbf{P}})=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is concerned, by the strong suppression of the weight $v_{\mathbf{P}}^{2}$ when the concentration decreases, in such a way that the identification of $n_{\rm CF}$ with $n_{\rm B}-n_{0}\simeq n_{\rm B}$ is always valid in the strong-coupling limit.
Finally, we discuss the profiles of the Fermi momentum distribution (39) for fixed coupling strength and different concentrations. In Fig. 6, $n_{\mathrm{F}}({\mathbf{k}})$ is displayed for $g=2.8$ and different values of $x$. For small $x$, most fermions remain unpaired and are described by a weakly renormalized Fermi step, while a small fraction of fermions bound in molecules are described by the small contribution $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$, with $n_{\rm CF}\ll n_{\rm F}$. As $x$ increases, the height of the Fermi step is progressively reduced, indicating the increasing importance of interaction effects. The position of the Fermi step, on the other hand, decreases only slightly from the non-interacting value $k_{\rm F}$. We notice that this behavior, together with the one just discussed for $n_{\rm CF}({\mathbf{P}})$ reflects again the breakdown of the Luttinger theorem that, in its modified version in the presence of a condensate [65], would require $k_{\rm F}^{2}=k_{\rm F,UF}^{2}+P_{\rm F,CF}^{2}$ (where $k_{\rm F,UF}$ and $P_{\rm F,CF}$ indicate the Fermi momenta for the unpaired and composite fermions, respectively).
@@
The insets of Fig. 6 finally compare the full strong-coupling expression (39) for $n_{\rm F}({\mathbf{k}})$ with its small $\Delta_{0}/E_{\rm F}$ expansion, Eqs.(41-43), which holds for $x\to 0$ and $x\to 1$, as mentioned above. The distribution function (41) resulting from this expansion (tot) is broken down in paired (CF) and unpaired (UF) contributions, described by $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$ and Eq. (43) for $n_{\rm UF}({\mathbf{k}})$, respectively. One sees from this comparison that the small $\Delta_{0}/E_{\rm F}$ expansion fully agrees with the full strong-coupling expression (39). One also sees how, for $x=1$, the unpaired contribution is already strongly suppressed for the value $g=2.8$ considered in Fig. 6, consistently with its vanishing contribution for $g\to\infty$ discussed above.

### v3

@@ 1 Introduction
Mixed systems of bosons and fermions are at the base of our understanding of the physical world. They feature in the Standard Model of elementary-particle physics as gauge and matter fields, respectively, in high-density quark matter as diquarks interacting with unpaired quarks [1], in nuclear matter in the interacting boson-fermion model for atomic nuclei with an odd number of protons or neutrons [2], and as boson-mediated interactions in soft [3] as well as condensed [4] matter. Ultra-cold atomic gases provide an exceptionally versatile platform for the simulation of quantum matter [5]. In the case of Bose-Fermi (BF) atomic mixtures, several realizations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have allowed exploration of a wide range of phenomena like phase separation [22], polarons [13, 17, 23, 26, 31], dual superfluidity [15, 16, 19, 20], collective excitations [34, 39], mediated interactions [28, 35, 36, 37], and Feshbach molecules [6, 8, 10, 25, 30, 33].
@@
In parallel, BF mixtures have also been realized in semiconductor nanostructures due to the recent advent of atomically thin transition metal dichalcogenides (TMD) [40, 41, 42, 43, 44, 45]. These materials intrinsically offer a remarkable opportunity to investigate the physics in two dimensions (2D) where topological phases play a crucial role [46, 47]. The so-called spin-valley locking, originating from the strong spin-orbit coupling of the transition metal atoms, allows for a complete distinguishability between electrons (and/or holes) populating the two spin-polarized valleys at Brillouin points $K$ and $K^{\prime}$ [48]. Exciton formation and charge doping are therefore two distinct processes that affect separate assemblies of electrons (and/or holes). In this scenario, the production of a BF mixture in a two-dimensional solid-state setting becomes therefore possible, with bosons corresponding to tightly bound excitons and fermions to doped charges. The recent detection of electrically tunable two-dimensional Feshbach resonances in TMD bilayers has put this target on more solid ground [43, 49].
@@
However, the unique flexibility of atomic gases is unrivaled when the need arises to manipulate dimensionality and interaction strength at the same time. Confinement-induced resonances accomplish this purpose by allowing for the realization of a dimensional crossover, starting from 3D confinement down to (quasi-)2D one. If the energy level spacing of the 3D trapping system is kept much larger than the energy scale of the intervening scattering processes, the system becomes effectively 2D and the ensuing 2D scattering length can be manipulated by varying either the scattering length associated with the original 3D Feshbach resonance and/or the confinement length, yet ensuring that the quasi-2D condition above remains satisfied [50, 51, 52]. This situation is particularly apt for BF mixtures, for which one would need to adjust independently the BF and the boson-boson (BB) interactions owing to the mechanical stability of the assembly being conditioned primarily by that of the Bose component [53]. Having at one’s disposal two independent degrees of freedom in the 3D scattering length and the confinement length, one can then correspondingly tune the effective BF and BB interactions (even though in a highly non-linear fashion [50]) and simultaneously analyze the stability of different phases of the mixture across the whole resonant regime.
@@
Within the above scenario, it is crucial to explore BF mixtures with a tunable BF attraction in a 2D setting. Until now, theoretical efforts have concentrated mainly on 3D settings [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. In 2D, theoretical work has been sparse so far [87, 88, 89, 90] and has primarily dealt with phase competition in optical lattices [89, 87], phase stability in atomic traps [88], or with the bosonic impurity limit of a boson density ${n_{\mathrm{B}}}$ much smaller than the fermion density ${n_{\mathrm{F}}}$ [90].
In this work, we study a 2D BF mixture in homogeneous space with boson and fermion densities spanning from the bosonic impurity limit (${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$) to the matched density regime (${n_{\mathrm{B}}}={n_{\mathrm{F}}}$). Similarly to the analyses carried out in 3D [76, 80, 84, 83, 86, 91], we are here interested in the competition between boson condensation and BF pairing in mixtures with an attractive and tunable BF interaction at zero temperature. For this reason, we consider concentrations of bosons $x={n_{\mathrm{B}}}/{n_{\mathrm{F}}}\leq 1$, such that a full competition between pairing and condensation is allowed. The BB interaction (when considered) is taken as repulsive for mechanical stability reasons. Numerical calculations focus on the case of equal boson and fermion masses in order to reduce the computational effort. However, to facilitate possible future studies, the theoretical formalism is developed for generic boson and fermion masses $m_{\rm B}$ and $m_{\rm F}$.
Our main results are as follows. (i) We find that the condensate is progressively depleted as the BF attraction increases, in analogy to the 3D case. However, while in 3D the condensate vanishes beyond a critical coupling strength, in 2D the condensate does not exactly vanish but rather becomes exponentially small with the coupling strength. (ii) Similarly to the 3D case, we uncover a nearly universal behavior of the condensate fraction, the bosonic momentum distribution, and Tan’s contact parameter with respect to the boson concentration. (iii) BF pairs form for sufficiently strong BF attraction. Quite generally, these composite fermions are hybridized states between a pure molecular bound state and an unpaired state made of a fermionic atom and a boson belonging to the condensate. We find that these hybridized states produce quasi-particles with rather peculiar Fermi liquid features. (iv) The boson momentum distribution function shows quite a rich behavior when the BF coupling strength and boson concentration are varied, which could in principle be tested in future experiments with BF mixtures.
@@ 2.1 The system Hamiltonian
On physical grounds, the following picture is expected when the BF coupling strength is varied. In the weak-coupling limit, bosons are expected to be almost fully condensed, with a small depletion of the condensate density $n_{0}$ determined just by the weak BB repulsion [94] for $g\to-\infty$. Fermions instead fill up a Fermi sphere. In the strong-coupling limit, bosons are instead expected to be expelled out of the condensate and form molecular fermions that fill up a Fermi sphere. We will discuss below to what extent, within the present approach, this simple picture continues to hold also in two dimensions.
@@ 2.2 Many-body $T$-matrix
A warning is however in order. In the limiting case of a single impurity in a Fermi gas, the present approach it has been shown to miss the polaron-to-molecule transition in 2D [97], which is found instead by more sophisticated approaches [98, 99]. We should thus bear in mind that such a shortcoming of the present approximation, which is due to an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory, may also extend at finite boson concentration $x$, in particular in the limit $x\to 0$ that approaches the single impurity case.
@@
Let us now see the rationale behind the choice of diagrams in [91], which we wish to extend to 2D. Since the pioneering work by Nozières and Schmitt-Rink [100] on the related problem of the BCS-BEC crossover in two-component Fermi gases, it is well known that ladder diagrams are able to capture pairing (molecular) correlations in the normal phase (see also [101] for a review). Only after the inclusion of this class of diagrams does the superfluid critical temperature recover the Bose-Einstein condensation temperature in the strong-coupling limit of the BCS-BEC crossover. In addition, for the same contact potential we are considering, ladder diagrams provide the leading self-energy in the weak-coupling limit of the BCS-BEC crossover [102]. They thus provide a sensible scheme to describe the whole BCS-BEC crossover, even in the intermediate coupling region in which fully controlled approximations are not available.
@@
whereby the $\Theta$ functions appearing in the integrand correspond to the Fermi and Bose distributions in the zero temperature limit. In particular, for the concentrations $x\leq 1$ and the range of BF and BB coupling strengths considered in this work, ${\mu_{\mathrm{B}}}$ is always negative, and hence $\Theta(-\xi_{\mathbf{k}}^{\mathrm{B}})=0$.
@@ 2.3 Boson and fermion self-energies
The bosonic self-energy is made of two terms: originating, respectively, from BF and BB interactions. For the former, we proceed as in the corresponding problem in the BCS-BEC crossover: the self-energy of one species is obtained by closing the $T$-matrix by a propagator of the other species.
@@ 2.4 Green’s functions, momentum distributions, and densities
Finally, due to the presence of a bosonic component at zero temperature, the Hugenholtz-Pines condition ([107, 108, 109]) for the corresponding chemical potential is imposed as
@@ 3.2 Fermionic self-energy and momentum distribution
The frequency integration in Eq. (12) is computed through a contour integration on the left-hand side of the complex plane, due to the presence of the convergence factor $e^{i\omega 0^{+}}$. The pole originating from $G_{\mathrm{B}}^{0}({\mathbf{P}}-{\mathbf{k}},\Omega-\omega)$ is located on the right-hand side of the complex plane because $\mu_{\rm B}<0$, and does not contribute to the integral. The contributions from the poles of $T_{\mathrm{SC}}$ yield instead
@@ 3.4 Complete system of semi-analytic equations in closed form and their solutions
The momentum distribution of the composite fermions $n_{\mathrm{CF}}(\mathrm{P})=v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is shown in Fig. 5 for a number of concentrations corresponding to those of Fig. 4 along with the case $x=1$. In the latter case, the mixture is essentially made of a single-component gas of non-interacting point-like fermionic molecules, as illustrated by the Fermi step behavior with occupation number approaching unity. As $x$ decreases, interaction effects, originating from the hybridization of the molecular and atomic states, become significant and generate a rather peculiar behavior of the momentum distribution function, as evidenced by its upward bending with increasing momentum. (A similar behavior was also found in [71] for a BF mixture in a 1D lattice.) Moreover, we notice that the Fermi step, in particular for small concentrations, is located well above the expected position if one were to assume the validity of the Luttinger theorem for the composite fermions.
@@
The Luttinger theorem states that the volume of the Fermi sphere of a Fermi liquid does not depend on interaction, and thus coincides with the volume of the Fermi sphere of the corresponding non-interacting Fermi gas [115]. Its validity has been extended to density-imbalanced two-component Fermi gases [116] and, within a two-channel model, to BF mixtures [65]. In both cases, it was shown that, provided the system is in the normal phase, the volumes of the Fermi spheres are separately conserved for the two Fermi components (in the BF mixture case, the two components correspond to renormalized fermionic atoms and molecules). This is indeed what is found in the strong-coupling limit of a 3D BF mixture in [84], since in this limit the condensate vanishes identically. In contrast, in the present 2D case the condensate density never identically vanishes, and in particular the hybridization energy $\Delta_{0}$ remains finite even for $g\to\infty$ (except in the trivial case $x=0$ and in the more interesting case $x=1$).
@@
In a BF mixture with a condensate, it was shown in [65] that it is the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules that should be conserved. Fig. 5 shows a clear violation of the latter version of the Luttinger theorem since the molecular Fermi momentum is always larger than $k_{\mathrm{F}}$, which is the Fermi momentum that should determine the total volume (i.e the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules). We attribute this shortcoming to two factors. First, the large values of the hybridization energy-scale $\Delta_{0}$ obtained in our theory, which are in turn caused by an overestimation of the atom-molecule repulsion (see Sec. 3.1), eventually lead to an overpopulation of the condensate. So, the simple strong-coupling picture of a Fermi-Fermi mixture of unpaired fermions and molecules obeying the Luttinger theorem is never reached in 2D. Second, as stressed in [116] and explicitly shown in [117], when dealing with approximate theories the validity of the Luttinger theorem is not guaranteed for schemes, like the present one, that do not use fully self-consistent fermionic Green’s functions in their formulation.
@@
Nevertheless, the breakdown of the Luttinger theorem in 2D for the hybridized composite fermions is compensated, as far as the integrated quantity $n_{\mathrm{CF}}=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}n_{\rm CF}({\mathbf{P}})=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is concerned, by the strong suppression of the weight $v_{\mathbf{P}}^{2}$ when the concentration decreases, so that the identification of $n_{\rm CF}$ with $n_{\rm B}-n_{0}\simeq n_{\rm B}$ is always valid in the strong-coupling limit.
Finally, we discuss the profiles of the Fermi momentum distribution (39) for fixed coupling strength and different concentrations. In Fig. 6, $n_{\mathrm{F}}({\mathbf{k}})$ is displayed for $g=2.8$ and different values of $x$. For small $x$, most fermions remain unpaired and are described by a weakly renormalized Fermi step, while a small fraction of fermions bound in molecules are described by the small contribution $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$, with $n_{\rm CF}\ll n_{\rm F}$. As $x$ increases, the height of the Fermi step progressively reduces, indicating the increasing importance of interaction effects. The position of the Fermi step, on the other hand, decreases only slightly from the non-interacting value $k_{\rm F}$. We notice that this behavior, together with the one just discussed for $n_{\rm CF}({\mathbf{P}})$ reflects again the breakdown of the Luttinger theorem that, in its modified version in the presence of a condensate [65], would require $k_{\rm F}^{2}=k_{\rm F,UF}^{2}+P_{\rm F,CF}^{2}$ (where $k_{\rm F,UF}$ and $P_{\rm F,CF}$ indicate the Fermi momenta for the unpaired and composite fermions, respectively).
@@
The insets of Fig. 6 finally compare the full strong-coupling expression (39) for $n_{\rm F}({\mathbf{k}})$ with its small $\Delta_{0}/E_{\rm F}$ expansion, Eqs.(41-43), which holds for $x\to 0$ and $x\to 1$, as mentioned above. The distribution function (41) resulting from this expansion (tot) is broken down into paired (CF) and unpaired (UF) contributions, described by $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$ and Eq. (43) for $n_{\rm UF}({\mathbf{k}})$, respectively. From this comparison, one sees that the small $\Delta_{0}/E_{\rm F}$ expansion fully agrees with the full strong-coupling expression (39). One also sees how, for $x=1$, the unpaired contribution is already strongly suppressed for the value $g=2.8$ considered in Fig. 6, consistently with its vanishing contribution for $g\to\infty$ discussed above.